### v1

# Numerical optimization for the compatibility constant of the lasso

Abstract
Compatibility condition and compatibility constant have been commonly used to evaluate the prediction error of the lasso when the number of variables exceeds the number of observations. However, the computation of the compatibility constant is generally difficult because it is a complicated nonlinear optimization problem. In this study, we present a numerical approach to compute the compatibility constant when the zero/nonzero pattern of true regression coefficients is given. We show that the optimization problem reduces to a quadratic program (QP) once the signs of the nonzero coefficients are specified. In this case, the compatibility constant can be obtained by solving QPs for all possible sign combinations. We also formulate a mixed-integer quadratic programming (MIQP) approach that can be applied when the number of true nonzero coefficients is moderately large. We investigate the finite-sample behavior of the compatibility constant for simulated data under a wide variety of parameter settings and compare the mean squared error with its theoretical error bound based on the compatibility constant. The behavior of the compatibility constant in finite samples is also investigated through a real data analysis.
Keywords: Lasso, Compatibility condition, Sparse estimation, Optimization
1 Introduction
The success of the lasso (least absolute shrinkage and selection operator; tibshirani1996lasso) lies in the computational efficiency (efron2004least; friedman2010regularization; boyd2011distributed; boyd2004convex; Stellato2020OSQP) and strong theoretical foundations in high-dimensional settings (bickel2009simultaneous; van2009conditions; buhlmann2011statistics; zhang2010nearly; zhao2006model; wainwright2009sharp; meinshausen2009lasso). In particular, the theory of the lasso has been shown to provide good prediction accuracy even when the number of variables exceeds the number of observations in practice. For example, some strong conditions, such as the compatibility condition, restricted eigenvalue condition, and irrepresentable condition, are sufficient conditions for obtaining a desirable property, such as oracle inequality and model selection consistency (zhao2006model; bickel2009simultaneous; van2009conditions; buhlmann2011statistics). The theory of the lasso shares many similarities with that of compressive sensing, where the restricted isometry property has been shown to guarantee sparse recovery (candes2005decoding; donoho2006compressed).
This paper focuses on the compatibility condition (buhlmann2011statistics), one of the crucial conditions to investigate the oracle inequality of the lasso estimator. The compatibility condition can be weaker than other assumptions, including the restricted eigenvalue, irrepresentable, and restricted isometry conditions (van2009conditions; bickel2009simultaneous; buhlmann2011statistics). Therefore, it is crucial to investigate whether the compatibility condition holds; if it does not, there is a possibility that the lasso can fail to achieve good prediction accuracy.
The compatibility condition provides an upper bound of the mean squared error (MSE) in a high-dimensional setting. This upper bound is highly dependent on the compatibility constant, say $\phi$. Therefore, the computation of the compatibility constant is essential in understanding the prediction error of the lasso. Indeed, as the compatibility constant becomes large, the theoretical upper bound decreases, resulting in improvement of prediction accuracy (buhlmann2011statistics). Nevertheless, little attention has been paid to evaluating the compatibility constant for a given design matrix. One of the main reasons would be that most theoretical research has focused on establishing upper bounds for evaluating the prediction error, and the compatibility constant is typically used only to provide a sufficient condition for guaranteeing such upper bounds.
However, understanding the finite-sample behavior of the compatibility constant would be essential for deepening the theoretical understanding of the lasso. For example, computing the compatibility constant enables us to investigate how large it is for a given design matrix, how it behaves with different numbers of observations and different numbers of dimensions, or how far the upper bound deviates from the prediction error. Such empirical investigations would provide a valuable step toward bridging the gap between the lasso theory and its practical performance.
In practice, the computation of $\phi$ is generally impossible as it depends on the zero/nonzero pattern of true regression coefficients that cannot be known in advance. If we assume that we know them in advance, the compatibility constant can be computed, but it can still be challenging in practice. The computation of the compatibility constant is similar to finding the smallest eigenvalue of the sample covariance matrix, but can differ and may be larger than the smallest eigenvalue as the corresponding vector (although not exactly an eigenvector) must lie within a nonconvex set defined by indices of true nonzero coefficients (buhlmann2011statistics). Therefore, it would be essential to develop a numerical approach to evaluate the compatibility constant in practice.
In this study, we present a numerical approach to compute the compatibility constant when the zero/nonzero pattern of true regression coefficients is known. We show that when the signs of the nonzero coefficients are provided, the computation reduces to a Quadratic Programming (QP). Using this property, we can obtain the compatibility constant by solving QP for all combinations of signs and selecting the minimum value. When the number of true nonzero coefficients is moderately large, it is difficult to obtain the compatibility constant with the above QP-based procedure due to a huge number of combinations of the signs of nonzero regression coefficients. To handle the combinatorial complexity, we formulate the problem as a mixed-integer quadratic programming (MIQP, BertsimasTsitsiklis1997; BertsimasWeismantel2005) incorporating Big-$M$ constraints (NemhauserWolsey1988) or special ordered set type 1 (SOS1) variables (BealeTomlin1970; Williams2013).
We investigate the numerical behavior of the compatibility constant for simulated data for a wide variety of parameter settings. The results show that the compatibility constant is highly dependent on the number of observations; therefore, we must take it into account when investigating the prediction error. Furthermore, we provide a heuristic approach to estimate the compatibility constant when the zero/nonzero pattern of true regression coefficients is unknown. This approach is applied to the analysis of the S&P500 dataset to demonstrate the relevance to the simulation results.
The remainder of this paper is organized as follows. In Section 2, we briefly review the compatibility condition and its related topics. In Section 3, we provide a method for computing the compatibility constant based on QP and MIQP. Section 4 presents numerical simulations to compute the compatibility constant in a wide variety of settings, and compares the MSE and its upper bound constructed with the compatibility constant. In Section 5, we provide an actual data analysis to illustrate the behavior of the estimate of the compatibility constant. Section 6 concludes our work.
2 Preliminaries
Let $\bm{y}=(y_{1},\dots,y_{n})^{\top}\in\mathbb{R}^{n}$ be a response vector and $X\in\mathbb{R}^{n\times p}$ be a design matrix. Suppose that predictors are scaled and responses are centered; that is, without loss of generality, the intercept is 0 and ${\rm diag}(X^{\top}X)=nI_{p}$, where $I_{p}$ is a $p\times p$ identity matrix. Consider the linear regression model
$\bm{y}=X\bm{\beta}+\bm{\varepsilon},$ (1)
where $\bm{\beta}$ is a regression coefficient vector and $\bm{\varepsilon}$ is an error vector. Consider the lasso estimation,
$\frac{1}{2n}\|\bm{y}-X\bm{\beta}\|_{2}^{2}+\lambda\|\bm{\beta}\|_{1},$
where $\lambda>0$ is a regularization parameter which controls the fit to the data and the degree of sparsity.
For constructing the theory of the lasso for high-dimensional data, we assume that the true regression coefficients are sufficiently sparse. Let $S:=\{j\mid\beta_{j}\neq 0\}$, and let $s:=|S|$ denote its cardinality, which is assumed to be sufficiently small compared with $n$ and $p$. We may also consider the case $n<p$, but this is not necessary; we assume that $\frac{\log p}{n}\rightarrow 0$ as $n\rightarrow\infty$.
The compatibility condition is one of the most important assumptions for deriving an upper bound on the mean squared error (MSE) in high-dimensional settings. Let $\hat{\Sigma}=X^{\top}X/n$ be the Gram matrix of the design matrix. The compatibility condition, a condition for the design matrix $X$, is defined as follows:
Definition 1.
The compatibility condition (buhlmann2011statistics, p. 106) is met if there exists some $\phi>0$ such that, for any $\bm{v}\in\mathbb{R}^{p}$ satisfying $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$, it holds that
$\|\bm{v}_{S}\|_{1}^{2}\leq s\frac{\bm{v}^{\top}\hat{\Sigma}\bm{v}}{\phi^{2}}$ (2)
Here, $\phi>0$ is referred to as the compatibility constant.
The compatibility constant is related to the smallest eigenvalue of $\hat{\bm{\Sigma}}$ because $\|\bm{v}_{S}\|_{1}^{2}\leq s\|\bm{v}_{S}\|_{2}^{2}$. However, $\phi$ is generally larger than the smallest eigenvalue because of the constraint $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$ (buhlmann2011statistics). The constraint $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$ restricts the vectors to be in a nonconvex set where the main direction of the vector mostly aligns with $S$. In other words, feasible directions cannot deviate too far from the coordinates indexed by $S$.
To illustrate this intuition, consider the compound-symmetry structure, $\hat{\Sigma}=\rho\bm{1}_{p}\bm{1}_{p}^{\top}+(1-\rho)\bm{I}_{p}$ with $\rho$ being the correlation coefficient between any two predictors. When $\rho>0$, the largest eigenvalue of $\hat{\Sigma}$ is $1+(p-1)\rho$ and the corresponding eigenvector is $\bm{1}_{p}$. The other eigenvalues are all $1-\rho$ and the corresponding eigenvectors are those orthogonal to $\bm{1}_{p}$. However, a direction $\bm{1}_{p}$ cannot satisfy the constraint $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$ when $|S^{c}|\gg|S|$. Consequently, the feasible region excludes the direction associated with the largest eigenvalue, and any feasible $\bm{v}$ must deviate substantially from the direction $\bm{1}_{p}$. Therefore, when $\rho\approx 1$, $\bm{v}^{\top}\hat{\Sigma}\bm{v}$ that satisfies $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$ must be small, leading to a small compatibility constant. In contrast, when $\rho\approx 0$, the quadratic form $\bm{v}^{\top}\hat{\Sigma}\bm{v}$ does not vary significantly across different directions $\bm{v}$. As a result, the compatibility constant tends to be larger when the predictors are weakly correlated. In other words, the compatibility condition may hold when the predictors are not highly correlated.
Under the compatibility condition, we obtain the following result (Tibshirani2017Sparsity).
Proposition 1.
Suppose that $\lambda$ satisfies $\lambda\geq 2\sigma\sqrt{2/n(1+\log\left(\frac{p}{\delta}\right))}$ with $\delta\in(0,1)$. Then, with probability at least $1-\delta$, we have
$\frac{1}{n}\|X\hat{\bm{\beta}}-X\bm{\beta}\|_{2}^{2}\leq\frac{9s\lambda^{2}}{\phi^{2}}.$ (3)
In particular, plugging $\lambda=2\sigma\sqrt{2/n(1+\log\left(\frac{p}{\delta}\right))}$ into the above equation, we have
$\frac{1}{n}\|X\hat{\bm{\beta}}-X\bm{\beta}\|_{2}^{2}\leq\frac{72\sigma^{2}s(1+\log(p/\delta))}{n\phi^{2}}$ (4)
For details of the result and its proof, please refer to Section 4.2 of Tibshirani2017Sparsity. We may obtain results whose constant value is slightly different from (3) (e.g., 9 might be changed to 4; buhlmann2011statistics), but the rate of convergence remains unchanged; roughly speaking, the error converges to 0 with high probability when $\log p/n\rightarrow 0$ as $n\rightarrow\infty$.
The compatibility constant $\phi$ plays an important role in evaluating the error bound. Since the upper bound in Proposition 4 is proportional to $\phi^{-2}$, the magnitude of the compatibility constant $\phi$ directly determines how large the bound of the MSE is. When $\phi$ is large, the upper bound becomes small, which implies that the lasso may achieve a meaningful prediction accuracy if the tuning parameter is properly chosen.
Several conditions other than the compatibility condition have been proposed for establishing theoretical properties of the lasso (van2009conditions). In particular, the irrepresentable condition and restricted eigenvalue condition are two popular conditions. Their purposes differ depending on the aspect of the problem. The irrepresentable condition is closely related to model selection consistency (zhao2006model), while the restricted eigenvalue condition and the restricted isometry property are often used to analyze prediction error (bickel2009simultaneous; candes2005decoding). On the other hand, the compatibility condition is related to oracle inequalities as described in Proposition 4. Among these conditions, the compatibility condition is known to be the weakest condition that still guarantees oracle inequalities. In fact, van2009conditions show that the compatibility condition is implied by both the restricted eigenvalue condition and the restricted isometry property. For this reason, the compatibility condition can be considered a flexible assumption for studying the prediction performance of the lasso in high-dimensional settings.
3 Numerical optimization for compatibility condition
As the analytical derivation of the compatibility constant is not easy due to the constraint $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$, we provide a numerical optimization method. First, we reformulate the compatibility condition. The compatibility constant in (2) can be computed with the following optimization problem:
$\phi^{2}=\min_{\bm{v}}\left\{\frac{s\bm{v}^{\top}\hat{\Sigma}\bm{v}}{\|\bm{v}_{S}\|_{1}^{2}}\ \middle|\;\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}\right\}.$ (5)
The above problem turns out to be scale-invariant; define $\bm{v}^{\prime}=a\bm{v}$ with any nonzero value $a$. Then, we observe that $\bm{v}^{\prime}$ and $\bm{v}$ provide the same compatibility constant. Therefore, without loss of generality, we may restrict attention to vectors satisfying $\|\bm{v}_{S}\|_{1}=1$, in which case the constraint $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$ reduces to $\|\bm{v}_{S^{c}}\|_{1}\leq 3$. Hence, the problem (5) can be reformulated as
$\phi^{2}=\min_{\bm{v}}\left\{s\bm{v}^{\top}\hat{\Sigma}\bm{v}\ \middle|\;\|\bm{v}_{S}\|_{1}=1,\ \|\bm{v}_{S^{c}}\|_{1}\leq 3\right\}.$ (6)
The above problem appears to be a quadratic programming (QP). However, the constraint $\|\bm{v}_{S}\|_{1}=1$ is not a convex set. As a result, it would be difficult to directly apply the QP or other convex optimization tools.
3.1 Quadratic Programming
To address the above-mentioned issue, we consider a much simpler problem where the sign of each element of $\bm{v}_{S}$, say $\bm{z}\in\{\pm 1\}^{s}$, is fixed. In this case, the constraint, $\|\bm{v}_{S}\|_{1}=1$ is expressed as
$\sum_{j\in S}z_{j}v_{j}=1$ (7)
with a constraint that $z_{j}v_{j}\geq 0\ (\forall j\in S)$. Since
$\left\{\bm{v}\in\mathbb{R}^{p}\middle|\;\sum_{j\in S}z_{j}v_{j}=1,\ z_{j}v_{j}\geq 0\ (\forall j\in S)\right\}$
is a convex set, we can easily obtain $\phi^{2}$ for fixed $\bm{z}$, say $\phi^{2}_{\bm{z}}$, by solving the following problem:
$\phi^{2}_{\bm{z}}$= $\min_{\bm{v}\in\mathbb{R}^{p},\;\bm{u}\in\mathbb{R}^{|S^{c}|}}s\bm{v}^{\top}\hat{\Sigma}\bm{v}$ (8)
subject to $\sum_{j\in S}z_{j}\,v_{j}=1,$
$z_{j}\,v_{j}\geq 0$ $(\forall j\in S),$
$-u_{j}\leq v_{j}\leq u_{j}\quad$ $(\forall j\in S^{c}),$
$u_{j}\geq 0$ $(\forall j\in S^{c}),$
$\sum_{j\in S^{c}}u_{j}\leq 3.$
All constraints in (8) are linear and the feasible region is convex. The objective function $\bm{v}^{\top}\hat{\Sigma}\bm{v}$ is also convex because $\hat{\Sigma}$ is positive semidefinite. Hence, for a fixed sign pattern, the optimization problem (8) becomes a standard convex quadratic program (QP). We can obtain $\phi^{2}_{\bm{z}}$ with a solver for quadratic programming, such as the OSQP library (Stellato2020OSQP). The compatibility constant is obtained by solving (8) for all possible sign patterns of $\bm{z}$ as below:
$\phi^{2}=\min_{\,z_{j}\in\{\pm 1\}^{s}}\ \phi_{\bm{z}}^{2}.$ (9)
Remark 1.
The compatibility constant is invariant under a global sign flip of $\bm{v}$. Indeed, if $\bm{v}$ satisfies the condition in (6), $\bm{v}$ also satisfies both $\|\bm{v}_{S}\|_{1}=1$ and $\|\bm{v}_{S^{c}}\|_{1}\leq 3$, and the objective value is unchanged because $\bm{v}^{\top}\hat{\Sigma}\bm{v}=(-\bm{v})^{\top}\hat{\Sigma}(-\bm{v})$. Therefore, when enumerating sign patterns of $\bm{v}_{S}$, we only need to search $2^{s-1}$ combinations of signs.
3.2 Mixed-Integer Quadratic Programming
The problem (9) can be easily solved when the number of elements in active set, $s$, is sufficiently small. However, since the number of sign patterns to be (9) is $2^{s-1}$ and grows exponentially, the computation of (9) becomes infeasible when $s$ is large.
To avoid enumerating all $2^{s-1}$ sign patterns explicitly, we formulate the problem as a mixed-integer quadratic program (MIQP, BertsimasTsitsiklis1997; BertsimasWeismantel2005). For each $j\in S$, we introduce nonnegative variables $v_{j}^{+},v_{j}^{-}\geq 0$ and a binary variable $b_{j}\in\{0,1\}$ to express $v_{j}$ as follows:
$v_{j}=v_{j}^{+}-v_{j}^{-},\quad v_{j}^{+}\leq Mb_{j},\quad v_{j}^{-}\leq M(1-b_{j}),$ (10)
where $M>0$ is a sufficiently large constant. This formulation is referred to as Big-$M$ formulation (NemhauserWolsey1988). With this formulation, the constraint $\|\bm{v}_{S}\|_{1}=1$ can be written as a linear equality
$\sum_{j\in S}(v_{j}^{+}+v_{j}^{-})=1,$
and the problem (6) reduces to the following mixed-integer quadratic program (MIQP):
$\min_{\begin{subarray}{c}\bm{v}_{S}^{+},\bm{v}_{S}^{-}\geq 0,\\ \bm{v}_{S^{c}}\in\mathbb{R}^{|S^{c}|},\;\bm{u}\geq 0,\\ \bm{b}\in\{0,1\}^{s}\end{subarray}}$ $\bm{v}^{\top}\hat{\Sigma}\bm{v}\;\mathrlap{\;\text{with }\bm{v}_{S}=\bm{v}_{S}^{+}-\bm{v}_{S}^{-}}$ (11)
subject to $\sum_{j\in S}\bigl(v_{j}^{+}+v_{j}^{-}\bigr)=1,$
$v_{j}^{+}\leq Mb_{j}$ $(j\in S),$
$v_{j}^{-}\leq M(1-b_{j})$ $(j\in S),$
$-u_{j}\leq v_{j}\leq u_{j}$ $(j\in S^{c}),$
$\sum_{j\in S^{c}}u_{j}\leq 3.$
Alternatively, we may replace the Big-$M$ constraints $v_{j}^{+}\leq Mb_{j}$ and $v_{j}^{-}\leq M(1-b_{j})$ in (11) with the SOS1 (special ordered set of type 1) constraint on $(v_{j}^{+},v_{j}^{-})$, which enforces that at most one of them is nonzero (BealeTomlin1970; Williams2013).
The mixed-integer problem can be handled by modern solvers such as Gurobi and is tractable even when $s$ is moderately large. In practice, however, the MIQP may require substantial computation time when $s$ is large. Therefore, we usually impose a time limit in our implementation so that the solver returns the best feasible solution within the time limit.
4 Monte Carlo Simulation
4.1 Evaluation of compatibility constant via QP
We investigate how the compatibility constant behaves and how tight the theoretical upper bound is under various sample sizes $n$ and dimensions $p$. Because the compatibility condition strongly depends on the correlation structure of the design matrix, we generate data with different correlation structures. The $p$-dimensional predictors are independently generated from $N_{p}(\bm{0},\bm{\Sigma})$, where $\bm{\Sigma}$ has the compound-symmetry structure ${\Sigma}=(1-\rho)I_{p}+\rho\bm{1}_{p}\bm{1}_{p}^{\top}$. The predictors are scaled to have mean zero and variances 1. The sample sizes are set to $n=100,200,300,400,500,750,1000,1500,2000$, and the numbers of variables are set to $p=20,50,100,200,500,1000,2000,5000$. The true coefficient vector $\bm{\beta}$ has $s=5$ nonzero elements, with indices $S\subseteq\{1,\dots,p\}$ chosen uniformly at random, and these nonzero elements are generated from $\mathrm{Unif}(1,2)$. The response variables are generated from the regression model in (1) with $\bm{\varepsilon}\sim N(\bm{0},\sigma^{2}I_{n}),$ where $\sigma^{2}$ is determined such that the signal-to-noise ratio (SNR) is equal to 1; that is, $\sigma^{2}:=\bm{\beta}^{\top}\bm{\Sigma}\bm{\beta}$. The regularization parameter is chosen as $\lambda=2\sigma\sqrt{\frac{2}{n}\left(1+\log\left(\frac{p}{\delta}\right)\right)}$ to have the bound of (4), where $\delta$ is set to 0.1.
For each iteration of the Monte Carlo simulation, we generate a dataset and compute the compatibility constant $\phi$ by solving the QP with the OSQP solver using the osqp package in R. The lasso estimator is implemented with glmnet. We obtain the MSE and its upper bound:
$\mathrm{MSE}=\frac{1}{n}\|\bm{X}(\hat{\bm{\beta}}-\bm{\beta})\|_{2}^{2},\qquad\mathrm{bound}=9\lambda^{2}s/\phi^{2}.$ (12)
Recall that the error bound in Proposition 4 is proportional to $1/\phi^{2}$. Therefore, when the compatibility constant is close to zero, the theoretical upper bound becomes extremely large and practically uninformative. The computation of $\phi$ under different correlation structures and sample sizes helps us understand how the prediction error and its upper bound based on $\phi$ behave in finite samples.
Figure 1: Compatibility constant $\phi$ for varying $(n,p,\rho)$. The red dashed line indicates the upper bound of compatibility constant for population covariance matrix, while the black dashed line indicates $\phi=0$.
Figure 1 presents the compatibility constant $\phi$ for different $n$, $p$, and $\rho$ under $R=10$ runs. The results show that the compatibility constant becomes large as the number of observations increases. This behavior can be explained by the fact that the rank of the Gram matrix $\hat{\Sigma}=\sum_{i=1}^{n}\bm{x}_{i}\bm{x}_{i}^{\top}/n$ is less than or equal to $\min(n,p)$. When $n$ is small, the Gram matrix $\hat{\Sigma}$ has low rank and a large null space, allowing vectors satisfying the condition $\mathcal{C}:=\{\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}\}$ to align with directions where the quadratic form $\bm{v}^{\top}\hat{\Sigma}\bm{v}$ becomes nearly zero. As $n$ increases, the rank of $\hat{\Sigma}$ increases and its null space shrinks, and as a result, it becomes harder to find such directions within $\mathcal{C}$.
We also observe that as $p$ increases, the compatibility constant decreases. This is probably because increasing $p$ enlarges the feasible set $\mathcal{C}$, which provides more degrees of freedom and makes it easier to find directions in which the quadratic form $\bm{v}^{\top}\hat{\Sigma}\bm{v}$ is small. In particular, when $p$ is large and $n$ is small, the compatibility constant sometimes becomes $\phi=0$, which implies the compatibility condition does not hold. Therefore, for high-dimensional data, several hundreds of sample sizes would be necessary to get a reasonable compatibility constant. The compatibility constant also highly depends on the value of $\rho$; as the value of $\rho$ increases, the value of $\phi$ decreases. This can be explained by noting that the compatibility constant behaves similarly to the smallest eigenvalue of the covariance matrix, $1-\rho$, which decreases as $\rho$ increases.
The red dashed line indicates the upper bound of compatibility constant for the population covariance matrix, $\Sigma$; when $s$ is even, $\phi^{2}=1-\rho$, and when $s$ is odd, $\phi^{2}\leq(1-\rho)\left(1+\frac{1}{s(p-s)}\right)$. For derivation of the above upper bound, please refer to Appendix A. We observe that when $p$ is large, the limit would converge to the true but slowly, so that we need a relatively large number of observations to obtain a reasonable bound derived from the compatibility constant of the population covariance matrix. In other words, the compatibility condition can fail even under moderate correlation structures, simply due to finite-sample limitations rather than the population correlation itself.
Figure 2: Scaled MSE and upper bound, ${\rm MSE}_{\rm scaled}$ and ${\rm bound}_{\rm scaled}$, for varying $(n,p,\rho)$.
Figure 2 shows the scaled mean squared error (MSE) and its upper bound as below
${\rm MSE}_{\rm scaled}=\frac{n}{\log p}{\rm MSE},\quad{\rm bound}_{\rm scaled}=\frac{n}{\log p}{\rm bound},$ (13)
where the MSE and its bound are defined in (12). Here, the “scaled” means multiplication of $n/\log p$ so that the right-hand side of inequality (4), ${\rm MSE}\lesssim C\log p/n$, has a quantity of order one.
The result shows that the scaled MSE is relatively stable regardless of the value of $n$ and $p$, except for small $n$. On the other hand, the upper bound is unstable when $n$ is small and $p$ is large. This is because the compatibility constant can be small for small $n$ as discussed in Figure 1. In particular, when $p$ is sufficiently large and $n$ is small, the upper bound cannot be obtained because of $\phi=0$. Therefore, the upper bound made by the small sample sizes can be much larger than that with large sample sizes.
Figure 3: Ratio between the MSE and the upper bound, $\text{MSE}/\text{bound}$, for varying $(n,p,\rho)$.
Figure 3 shows the ratio between the MSE and its upper bound. The ratio becomes extremely large for high-dimension and small sample sizes. In particular, the ratio can be $10^{5}$ to $10^{6}$, which implies the upper bound does not make sense to evaluate the MSE. Therefore, the compatibility constant cannot be useful unless $n$ is relatively large. For large sample sizes, the ratio is $10^{1}$ to $10^{2}$, which is reasonable to evaluate the MSE but still has a gap between MSE and bound. As a result, when the compatibility constant is sufficiently large, the MSE would be small so that the lasso might perform well. Even if the compatibility constant is small, there exists situations where the MSE is sufficiently small compared with its bound.
4.2 Computation of upper bound via MIQP
The QP with all sign combinations described above can be applied only when $s$ is sufficiently small, such as $s=5$. When $s$ is large, such as $s=50$, the QP approach becomes infeasible even with parallel computation. In this case, mixed-integer quadratic programming (MIQP) provides a more feasible alternative.
First, we consider a small problem for which the QP can be computed within a realistic time, and compare the results with those from MIQP. For MIQP, we employ warm starts and early stopping with a time limit to evaluate the upper bound of the compatibility constant, and investigate how accurate this upper bound is.
Figure 4 shows the compatibility constant obtained by MIQP and QP (upper panel) and their ratio, $\phi_{\mathrm{MIQP}}/\phi_{\mathrm{QP}}$ (lower panel). In the upper panel, the gray shaded regions summarize the distribution of the compatibility constant obtained by QP: the dark gray band represents between the first and third quartiles, while the light gray bands show the overall ranges from the minimum to the maximum. The boxplots show the MIQP-based estimates for different warm-start sizes $K$ and time limits. The warm start value is obtained by solving the QP with a fixed sign vector $\bm{z}$. A set of $K$ ($K=1,5,10,20$) random sign vectors $\bm{z}$ is generated from a Bernoulli distribution. For each vector, we solve the corresponding QP and select the one that yields the smallest objective value. The time limits for MIQP are set to $T=30,60,120,$ and $180$ seconds. The other parameters are set to $p=2000$ (i.e., relatively high-dimensional case), $\rho=0.4$, $n=500$, $1000$, $2000$, and $s=5$ and $10$.
Figure 4: Compatibility constant obtained by MIQP and QP (upper panel) and their ratio, $\phi_{\mathrm{MIQP}}/\phi_{\mathrm{QP}}$ (lower panel). In the upper panel, the gray shaded regions summarize the distribution of the compatibility constant obtained by QP: the dark gray band represents between the first and third quartiles, while the light gray bands show the overall ranges from the minimum to the maximum. The boxplots show the MIQP-based estimates for different warm-start sizes $K$ and time limits.
For all cases, the compatibility constant obtained by MIQP is always larger than that obtained by QP because MIQP uses a time limit and cannot always attain the optimal value. The difference between MIQP and QP is relatively small, especially for large $n$. When $n$ is small, MIQP can yield a relatively larger $\phi$ than QP, and thus the difference is not negligible. Nevertheless, the ratio between the two compatibility constants, $\phi_{\mathrm{MIQP}}/\phi_{\mathrm{QP}}$, never exceeds 2, implying that the order of the error bound may not change due to the difference between MIQP and QP.
The ratio remains almost the same across different time limits for a fixed $K$, indicating that the warm start provides a good initial value. In particular, when $K$ is large, the ratio becomes quite small. We also observe that the effect of $K$ is much larger than that of the time limit. This behavior suggests that most of the improvement can be almost achieved by the warm start based on $K$ random signs, and MIQP has only a limited search space left to explore. Therefore, increasing the time limit gives almost no additional benefit, while increasing $K$ significantly enhances the quality of the warm start. Thus, it is preferable to choose a large $K$, while the time limit does not necessarily have to be large; 60 seconds appears to be sufficient when $K$ is large.
Next, we consider a larger problem with $s=50$, where the QP with all sign combinations cannot be performed in practice due to the high computational cost of $2^{50-1}$ patterns. We compute an upper bound of $\phi$ using MIQP described above. Based on Figure 4, we set $K=20$ and a time limit of $T=60$ seconds. The other settings are the same as in Figure 1. The results are shown in Figure 5. The red dashed line indicates the compatibility constant for the population covariance matrix, $1-\rho$, while the black dashed line indicates $\phi=0$.
Figure 5: Compatibility constant $\phi$ with $s=50$ with for varying $(n,p,\rho)$. The red dashed line indicates the compatibility constant for the population covariance matrix, $\sqrt{1-\rho}$, while the black dashed line indicates $\phi=0$.
The results show similar tendency to Figure 1, but yields much lower $\phi$ for small sample sizes. In particular, when $p=5000$ and $n=1000$ (high-dimensional and relatively large sample sizes), the upper bound of compatibility constant is 0, indicating that the compatibility constant is also 0. Therefore, even when the number of observations is large, the compatibility constant can be zero or close to zero when $s$ is relatively large. This result shows a clear finite-sample gap compared with asymptotic intuition. More surprisingly, even in the uncorrelated case ($\rho=0$), the compatibility condition fails simply due to the relationship among $(n,p,s)$. Therefore, the finite-sample behavior is strongly dominated by these three factors rather than true correlation among predictors. The MIQP with time limit cannot always be exact, but is helpful here because it can detect that $\phi$ becomes zero or close to zero in a challenging settings. This observation cannot be obtained by QP alone due to its high computational cost.
5 Real data analysis
We investigate how the compatibility constant $\phi$ and the error bound behave as the sample size increases through the analysis of real data. We use S&P 500 constituents to investigate the relationship between the daily return of Apple and that of the remaining index members. Since the composition of the S&P 500 index changes over time, not all the 500 stocks are continuously observed over the entire period. We download data from Yahoo Finance for the period from November 2017 to October 2025, and retain only the stocks for which daily records are available throughout this period. This screening results in 476 stocks, one of which is Apple.
In the regression analysis, the daily return of Apple is treated as the response variable, while the returns of the remaining 475 stocks are used as predictors. We consider the arithmetic daily returns
$R_{t,j}=\frac{P_{t,j}-P_{t-1,j}}{P_{t-1,j}},$
where $P_{t,j}$ denotes the closing price of stock $j$ on day $t$. The data set consists of $n=2009$ observations and $p=475$ predictors. In what follows, the predictors are standardized and the response vector $\bm{y}$ is centered.
Figure 6: Histogram of correlations among predictors.
Figure 6 shows the histogram of correlations among predictors. Most pairs of variables are moderately and positively correlated, while a few exhibit relatively high correlations. This correlation structure may influence the value of the compatibility constant, although its effect appears to be limited.
To compute the compatibility constant, we need the zero/nonzero pattern of the true coefficients, which cannot be provided in practice. Therefore, we provide a heuristic method to estimate the zero/nonzero pattern of the true coefficients. First, we divide the dataset into training and test datasets. The number of observations for the training and test data is $n_{\rm train}=1009$ and $n_{\rm test}=1000$, respectively. Then, we fit the lasso and estimate the set $S$ as follows:
  1. 1.
    We first select the regularization parameter $\lambda$ by the 10-fold cross validation. The cross-validation is known to be inconsistent in model selection and tends to select overly complex models (shao1993linear). Indeed, the number of nonzero coefficients selected by the cross-validation is $77$, which is relatively large.
  2. 2.
    The fitted value from cross-validation is expected to be sufficiently small because cross-validation is an unbiased estimator of the prediction error. Therefore, the unbiased estimator of $\sigma^{2}$, $\hat{\sigma}^{2}=\frac{1}{n_{\mathrm{train}}-s_{\mathrm{CV}}-1}\bigl\|\bm{y}_{\mathrm{train}}-X_{\mathrm{train}}\hat{\bm{\beta}}^{(\mathrm{CV})}\bigr\|_{2}^{2},$ would be a good estimator. Numerically, it is $\hat{\sigma}^{2}=1.268\times 10^{-4}.$
  3. 3.
    Based on the above $\hat{\sigma}^{2}$, we estimate the penalty level based on the Proposition 4 by
    $\lambda_{\mathrm{train}}=2\hat{\sigma}\sqrt{\frac{2}{n_{\mathrm{train}}}\Bigl(1+\log\frac{p}{\delta}\Bigr)}$ (14)
    with $\delta=0.1$. We get $\lambda_{\mathrm{train}}=3.085\times 10^{-3}.$
  4. 4.
    We fit a lasso on $(X_{\mathrm{train}},\bm{y}_{\mathrm{train}})$ again but with the above $\lambda_{\mathrm{train}}$. With the estimated regression coefficients, say $\hat{\bm{\beta}}^{(\mathrm{train})}$, we get a set of nonzero coefficients, $\hat{S}=\operatorname{supp}(\hat{\bm{\beta}}^{(\mathrm{train})})$.
In our experiment, we obtain the following $|\hat{S}|=11$ companies: Microsoft Corp., Amazon.com Inc., Meta Platforms Inc., Berkshire Hathaway Inc. (Class B), Mastercard Inc., Cisco Systems Inc., Amphenol Corp., IDEXX Laboratories Inc., Fastenal Co., CDW Corp., and Skyworks Solutions Inc.
The estimate of the active set, $\hat{S}$, can be used to compute the compatibility constant with the test dataset. To investigate the behavior of the compatibility constant with various sample sizes, we use the first $n=100k$ ($k=1,2,\dots,10$) observations of the test block. Denote the resulting subsample by $(X_{n},y_{n})$. For each $n$, we compute the empirical compatibility constant of $X_{n}$ with respect to the fixed set $\hat{S}$, say $\phi_{n}=\phi\bigl(\hat{S};X_{n}\bigr)$, using OSQP.
We also estimate the MSE and its bound computed with the $\phi_{n}$ to obtain the results similar to the Figures 13 in the simulation study. For a given $n$, we define the regularization parameter as $\lambda_{n}=2\hat{\sigma}\sqrt{\frac{2}{n}\Bigl(1+\log\frac{p}{\delta}\Bigr)},$ with $\delta=0.1$, which is the analog of (14) with $n$ in place of $n_{\mathrm{train}}$. We then compute the lasso estimate, $\hat{\bm{\beta}}_{n}$, with the above regularization parameter. The MSE in (12) is then estimated as
$\widehat{\mathrm{MSE}}=\frac{1}{n}\bigl\|X_{n}(\hat{\bm{\beta}}_{n}-\hat{\bm{\beta}}^{\rm train})\bigr\|_{2}^{2},\quad\mathrm{Bound}_{n}:=\frac{9\,|\hat{S}|\,\lambda_{n}^{2}}{\phi_{n}^{2}}.$
Here, the true $\bm{\beta}$ and $S$ are replaced by their estimates with training data, $\hat{\bm{\beta}}^{\rm train}$ and $\hat{S}$, respectively.
Figure 7 shows the estimate of compatibility constant $\phi_{n}$, the MSE and its error bound scaled by $n/\log p$ as described in (13) in the simulation study, and their ratio, $\mathrm{Ratio}_{n}={\mathrm{Bound}_{n}}/\widehat{\mathrm{MSE}}$.
Figure 7: The estimate of compatibility constant $\phi_{n}$ (left panel), the MSE and its error bound scaled by $n/\log p$ as described in (13) in the simulation study (middle panel), and their ratio, $\mathrm{Ratio}_{n}={\mathrm{Bound}_{n}}/\widehat{\mathrm{MSE}}$ (right panel).
The compatibility constant is a monotone non-decreasing function of $n$, which implies a large sample size is required to achieve a good error bound. For $n=100$ and $n=200$, $\phi_{n}=0$ and then the compatibility condition may not hold. Meanwhile, for large $n$, the compatibility constant can be greater than $0.3$, but it is still not very large. Considering that there exist a few pairs of large correlations in predictors, as shown in Figure 6, the moderate value of $\phi_{n}$ seems reasonable.
The scaled $\widehat{\mathrm{MSE}}$ is almost constant as a function of $n$, while the bound is a monotone decreasing function, which aligns with the simulation results. This monotonicity is likely due to the fact that the compatibility constant is non-decreasing in $n$. The error bounds for $n=100$ and $200$ cannot be obtained because the compatibility condition does not hold. The ratio is almost a monotone decreasing function, which also align with the results from simulation study. The ratio is not small but not too large compared to the results from the simulation study (see Figure 3). As a result, the overall results almost align with the simulation study and may provide relatively good accuracy when a sufficient number of observations is available.
6 Concluding remarks
We have provided an approach to compute the compatibility constant when the true zero/nonzero pattern of the true regression coefficients is known. The computation of the compatibility constant reduces to a QP if the signs of the true nonzero regression coefficients are given. Thus, we can compute the compatibility constant by solving QP for all combinations of signs. When the number of nonzero coefficients $s$ is large, we may alternatively use the MIQP with Big-$M$ or SOS1 under a time limit to obtain an upper bound of the compatibility constant. A more efficient computation for large $s$ would be crucial, but we will leave this as a future research topic.
The compatibility condition is one of the most important conditions in the lasso literature for achieving good accuracy in high-dimensional settings. More generally, researchers are interested in the theoretical properties of statistical methods under certain conditions. These conditions must be realistic, but in many cases they cannot be evaluated using a given dataset. This study attempted to address this issue by evaluating a key regularity condition of the lasso, which may serve as a guideline for assessing whether the method is likely to perform well. We believe that methods for checking such conditions using a given dataset would be useful in many practical situations. In the future, we would like to develop numerical approaches to check the regularity conditions and assess whether a method can be expected to perform well.
Acknowledgments
This work was supported by JSPS KAKENHI Grants JP23K11007, JP23H00466, JP23K22410, and JP23K01333.
Appendix A Compatibility constant under compound symmetry
We study the compatibility constant for a covariance matrix that has a compound symmetry, expressed as
$\Sigma_{\rho}:=(1-\rho)I_{p}+\rho\mathbf{1}_{p}\mathbf{1}_{p}^{\top},\qquad 0\leq\rho<1.$
Here, the correlation coefficient $\rho$ can be negative value, but for simplicity, we only consider the case where $\rho\geq 0$. Recall that the compatibility constant is defined as (6). For any $\bm{v}\in\mathbb{R}^{p}$, we have
$\bm{v}^{\top}\Sigma_{\rho}\bm{v}=(1-\rho)\|\bm{v}\|_{2}^{2}+\rho\,(\mathbf{1}^{\top}\bm{v})^{2}=(1-\rho)\|\bm{v}\|_{2}^{2}+\rho\left(\sum_{j=1}^{p}v_{j}\right)^{2}.$ (15)
The compatibility constant is closely related to the smallest eigenvalue of the covariance matrix. The eigenstructure of $\Sigma_{\rho}$ is simple: there exists a unique maximum eigenvalue $\lambda_{\max}=1+(p-1)\rho$ with the eigenvector being $\mathbf{1}$, and for the other (p1)(p-1)-dimensional vectors that is orthogonal to $\bm{1}$, we have the same eigenvalues $\lambda_{\min}=1-\rho$. Thus, the quadratic form $\bm{v}^{\top}\Sigma_{\rho}\bm{v}$ is smallest when $\bm{v}$ lies close to the subspace orthogonal to $\mathbf{1}$, that is, when $\sum_{j=1}^{p}v_{j}\approx 0$. This is also reasonable from Eq. (15) as the second term penalizes the component along $\mathbf{1}$.
If we minimize $\bm{v}^{\top}\Sigma_{\rho}\bm{v}$ with $\|\bm{v}\|_{2}=1$, it would simply mean choosing $\bm{v}$ as an eigenvector corresponding to the smallest eigenvalue, $1-\rho$. However, the constraint on the compatibility constant uses $\|\bm{v}_{S}\|_{1}=1$ rather than $\|\bm{v}\|_{2}=1$. Therefore, the decomposition of the quadratic form in (15) implies that we want to minimize $\|\bm{v}\|_{2}^{2}$ under the $\ell_{1}$-constraint $\|\bm{v}_{S}\|_{1}=1$, while we also want $\sum_{j=1}^{p}v_{j}$ to be as close to 0 as possible to avoid the direction corresponding to the largest eigenvalue.
We first derive a lower bound of $\phi$. Eq. (15) is evaluated with
$s\bm{v}^{\top}\Sigma_{\rho}\bm{v}\geq s(1-\rho)\|\bm{v}\|_{2}^{2}\geq s(1-\rho)\|\bm{v}_{S}\|_{2}^{2}\geq(1-\rho)\|\bm{v}_{S}\|_{1}^{2}=(1-\rho).$ (16)
The last inequality is derived from Cauchy-Schwarz inequality with two vectors, $\bm{v}_{S}$ and ${\rm sign}(\bm{v}_{S})$. Eq. (16) implies that the lower bound of $\phi^{2}$ is $1-\rho$.
If the upper bound of $\phi^{2}$ is also $1-\rho$, the compatibility constant is $\phi^{2}=1-\rho$. Fortunately, it is true when $s$ is even. When $s$ is odd, the derivation of exact value of $\phi^{2}$ would be more challenging. Therefore, we obtain an upper bound that is slightly greater than $1-\rho$.
Proposition 2.
It is shown that
$\begin{aligned} \phi^{2}&=1-\rho&\quad&(s:\ \text{even}),\\ \phi^{2}&\leq(1-\rho)\left(1+\frac{1}{sr}\right)&\quad&(s:\ \text{odd}),\end{aligned}$ (17)
where r:=psr:=p-s.
Proof.
Without loss of generality, we can define $S=\{1,2,\dots,s\}$ and $S^{C}=\{s+1,\dots,p\}$. Firstly, we assume that $s$ is even. Define $\bm{v}$ by
$v_{j}=\begin{cases}+\frac{1}{s},&j=1,\dots,\frac{s}{2},\\[2.0pt] -\frac{1}{s},&j=\frac{s}{2}+1,\dots,s,\\[2.0pt] 0,&j\in S^{c}.\end{cases}$ (18)
The vector $\bm{v}$ satisfies the constraint in (6) because $\|\bm{v}_{S}\|_{1}=1$ and $\|\bm{v}_{S^{c}}\|_{1}=0\leq 3$. Moreover, since $\sum_{j=1}^{p}v_{j}=0$ and considering $\|\bm{v}\|_{2}^{2}=\frac{1}{s}$, we obtain
$s\,\bm{v}^{\top}\Sigma_{\rho}\bm{v}=1-\rho.$
Therefore, $\phi^{2}=1-\rho$ when $s$ is even.
Next, we derive an explicit upper bound of $\phi$ when $s$ is odd. It is difficult to define $\bm{v}$ as in (18) becuase $s$ is odd, but we may define $\bm{v}$ that is similar to that in the even case in (18):
$v_{j}=\begin{cases}+\frac{1}{s},&j=1,\dots,\frac{s-1}{2},\\[2.0pt] -\frac{1}{s},&j=\frac{s-1}{2}+1,\dots,s,\\[2.0pt] \frac{1}{sr},&j\in S^{c}.\end{cases}$ (19)
The vector $\bm{v}$ satisfies the constraint in (6) because $\|\bm{v}_{S}\|_{1}=1$ and $\|\bm{v}_{S^{c}}\|_{1}=1/s\leq 3$. Furthermore, $\bm{v}_{S^{c}}$ is selected so that $\sum_{j=1}^{p}v_{j}=0$ is satisfied. Since $\|\bm{v}_{S^{c}}\|_{2}^{2}=1/(s^{2}r)$, we have
$s\,\bm{v}^{\top}\Sigma_{\rho}\bm{v}=s(1-\rho)\left(\frac{1}{s}+\frac{1}{s^{2}r}\right)=(1-\rho)\left(1+\frac{1}{sr}\right),$
which is the upper bound for $\phi^{2}$